OpenStudy (quickstudent):

When finding the slope of a line, instead of using the slope formula, wouldn't it be much easier to simply choose a point on the line and divide the y-value by the x-value? It gives the same answer much faster and easier.

OpenStudy (anonymous):

no

OpenStudy (anonymous):

you always want to do y2-y1/x2-x1

OpenStudy (quickstudent):

It gives the same answer much faster and easier.

OpenStudy (texaschic101):

lets try your way......points are : (2,3) and (4,5) slope formula : (y2 - y1) / (x2 - x1) slope = (5 - 3) / (4 - 2) slope = 2/2 = 1 now pick one of those points....(2,3).....now divide y by x.....3/2 = 1.5...not the same...or use (4,5)....now divide 5/4 = 1.25..not the same

OpenStudy (quickstudent):

Oh, I get it. The problem was I was only thinking about lines that pass through the origin. Sorry.